# The SAGA Survey. III. A Census of 101 Satellite Systems around Milky Way-mass Galaxies

### v2

@@ 3.4 Combining Spectroscopic and Photometric Data Sets
As the photometric catalog has been updated a few times during the course of the SAGA Survey, we combine the spectroscopic and photometric measurements by matching the sky coordinates. For each spectroscopic measurement, we search the photometric catalog within a 20′′ radius. Given the SAGA Survey depth, the photometric sources are not too crowded and in most cases only one unique match exists. When multiple photometric sources are possible matches, we identify the best match based on the luminosities and sizes of those photometric objects. The specific order by which a photometric source is selected as the best match to a spectroscopic measurement is described in the list in the SAGA code base.
@@ 4.2 Derived Properties
Radius and surface brightness: For most galaxies, the circularized half-light radius is derived directly from the axis ratio and semi-major axis from the Legacy Surveys catalog measured in -band. However, for galaxies available in the Siena Galaxy Atlas, we use as the half-light radius (one-third of the isophotal radius at 26 mag arcsec-2). We also use GALFIT to re-fit the photometry for a handful of galaxies that clearly have a bad photometric fit in the original catalog by visual inspection. We then calculated the effective surface brightness with the following formula:
@@
Stellar mass: As in Paper II, we estimate stellar mass using the Bell et al. (2003) relation, but calibrated the relation to recent estimates from Zibetti et al. (2009) and Taylor et al. (2011). We also assume a Kroupa (2001) initial mass function and an absolute solar -band magnitude of 4.65 (Willmer, 2018). The resulting stellar mass conversion is:
@@
For SAGA satellite galaxies (Section 4.3), we also measure the following quantities.
@@
Star Formation Rate (SFR): we calculate the star formation rate based on the both H equivalent width () and NUV flux. See Section 2.4 of Paper IV for details.
@@
“Quenched” flag: we label a galaxy as quenched if it has no significant H emission () and if it has a specific SFR (sSFR) in NUV below . See Section 3.1 of Paper IV for more detail.
@@ 4.3 SAGA Satellites
As in Paper II, we define a “satellite” as a galaxy that satisfies the following criteria: (i) fainter than the host galaxy, (ii) within 300 kpc in projection to the host galaxy, and (iii) has a heliocentric velocity that is within 275 km s-1 of the host galaxy. From the SAGA redshift catalog around 101 hosts, we identified 378 satellites in total, with 229 whose redshifts were first obtained by the SAGA Survey. Galaxies that do not meet these criteria are considered to be “background” galaxies (a handful of these galaxies are actually foreground galaxies; however, we will use the term background galaxies to refer to any non satellites).
@@
Figure 5 demonstrates this definition, with the satellite galaxies being in the middle left part of the plot. It is important to note that this observationally motivated satellite definition is not an attempt to strictly select objects that lie within the three-dimensional physical virial radius of their dark matter host halo, nor to strictly select objects that are gravitationally bound to the host. With only light-of-sight velocities and sky positions available, we chose a definition that is straightforward to implement in both observational and simulation data.
@@
With simulations, we can estimate the fraction of interlopers (galaxies that sit outside the three-dimensional radius in the light-of-sight direction) and potential fly-by galaxies (galaxies that are with the three-dimensional radius but have a speed greater than the escape velocity; Sinha & Holley-Bockelmann, 2012; An et al., 2019). Based on gravity-only simulations, the SAGA satellite definition gives an average interloper fraction of 30% for MW-mass halos. This fraction varies with projected distance, and drops to 15% for galaxies within 150 kpc in projection, as demonstrated in Paper II. Note that the radius considered here is always fixed to be 300 kpc, even though the individual SAGA systems could have a different virial radii, which are not measurable observationally.
@@
As for potential fly-by galaxies, Figure 5 highlights the fact that a small fraction (4%) of SAGA satellites may be unbound, indicated by purple squares (satellites) sitting outside the red solid bands, which represent the circular velocity of an NFW profile with and (red solid lines) plus the typical spectroscopic velocity error of 60 km s-1 in our measurement. The fraction of unbound objects observed is consistent with the gravity-only simulations when we use the SAGA satellite definition to select the halos in the simulations.
@@
Figure 5 also shows the velocity dispersion measured from all SAGA satellites. Satellites are binned in projected distances. For each bin, we fit a single Gaussian distribution, with its mean and variance being the free parameters, convolved with a constant velocity measurement error of 60 km s-1. Because we use all satellites to measure the velocity dispersion profile, it represents the stacked profile where systems with more satellites are weighted more. The velocity dispersion declines with projected radius to the virial radius, suggesting we are measuring a true dynamical mass. The increase beyond 300 kpc is likely due to increased contribution from interlopers. We are not able to measure the velocity dispersion for individual systems due to the small number of satellites in each system.
@@ 4.4 SAGA Background Sample
This sample of low-redshift and low-mass background galaxies, although not the main focus of the SAGA Survey, has enabled a range of studies including optimizing machine learning algorithms to select low-mass galaxies (Wu et al., 2022; Darragh-Ford et al., 2023), calibrating photometric sources for low-mass galaxy weak lensing measurements (Thornton et al., 2023), and measuring the mass loading factor in low-mass galaxies (Kado-Fong et al., 2024).
@@ 4.5 Likely Satellite Candidates and Incompleteness Correction
The quenched labels for unconfirmed candidates are determined by the following color–magnitude relation.
@@ 5.2.1 Individual SAGA stellar mass functions
We first show the cumulative satellite stellar mass function (SMF) for each of the 101 SAGA systems in Figure 8. While it can be difficult to discern individual SMFs in this figure, the main purpose of this figure is to demonstrate the system-to-system spread of the SMFs. The inset of this figure also allows us to quickly see the distribution of the number of satellites for the Gold and GoldSilver sample. Note that the SMFs and the number of satellites shown here already include the incompleteness correction described in Section 4.5 and Section B.1. Four of the 101 systems have the same level of survey completeness but do not have any confirmed satellites; their incompleteness corrected SMFs are shown in this figure as well.
@@ 5.2.2 Comparison with MW and M31 SMFs
The satellite SMF of MW and M 31 are also shown in Figure 8, and they are both well within the system-to-system spread of the SAGA SMFs. In particular, the existence of an LMC-mass satellite is not uncommon; about one-third (34/101) of the SAGA systems have at least one confirmed satellite with a stellar mass above . In the inset, one can see that the number of satellites in M 31 is on the high end of the SAGA systems, which is expected given M 31’s higher mass compared to SAGA hosts (Figure 3). On the other hand, the number of satellites in the MW is just slightly below the median number of the SAGA systems for the GoldSilver sample. If one only considers the expected correlation between host galaxy stellar mass and number of satellites, the MW would appear to have a lower satellite number given that the majority of SAGA hosts are less massive than the MW.
Note that in the main panel, the low-mass end of the MW SMF is closer to SAGA SMFs that are colored in blue, while the high-mass end of the MW SMF is closer to SAGA SMFs that are colored in yellow. In other words, the number of satellites in the MW (to the SAGA mass limit; i.e., GoldSilver sample) is lower than that of the SAGA systems that also have massive satellites like the LMC. This observation suggests that the MW satellites do not follow the strong correlation between the number of satellites and the most massive satellite mass seen in SAGA; we will demonstrate this correlation quantitatively in Section 5.7. This discrepancy suggests that the MW may have an atypical formation history compared to the SAGA hosts, which we will discuss in detail in Section 6.1.
@@ 5.3 Satellite Quenched Fraction
As Figure 9 demonstrates, we expect that a subset of low-mass quenched satellites exist in these systems for which we do not have a confirmed redshift. Since we have a good estimate of the expected number of these unconfirmed satellites, we include them in the calculation of the quenched fraction. For unconfirmed satellites, the quenched definition is based on their color and absolute magnitude, as described in Eq. 6. Note that the incompleteness corrected and uncorrected results are very similar in the Gold sample (right of the vertical gold dashed line), and deviate from each other at lower masses.
The satellite quenched fraction in the 101 SAGA systems is marginally higher than, but statistically consistent with, what we reported in our Stage II results (Paper II). When compared with the satellite quenched fraction in the Local Group, SAGA systems’ average satellite quenched fraction is lower for stellar mass below . This difference between SAGA systems and Local Group is unlikely due to observational incompleteness for three reasons:
@@
First, if the SAGA Gold sample had an overall satellite quenched fraction of 70% like the Local Group, it would imply there were more than 200 quenched satellites with that were not identified by our survey. This is extremely unlikely given our completeness estimates, and would significantly alter the stellar mass function if it were true.
@@ 5.4 Satellite Properties
The color–mass relation (left panel) shows a visible separation between red quenched satellites and blue star-forming satellites, although some star-forming satellites have redder colors than expected for their mass. As a result, there is no simple color–mass cut that can cleanly separate quenched and star-forming satellites. The color–mass cut used by Carlsten et al. (2022) to separate quenched and star-forming satellites is shown as a thin brown line on the plot. Notably, the LMC and the SMC (the two pink stars with the highest stellar masses) are bluer than any of the SAGA satellites of similar mass, possibly as a result of star formation triggered by their ongoing infall into the MW (Shao et al., 2018). We refer readers to Section 7.2 of Paper IV for more discussion.
@@ 5.5 Satellite Radial Distribution
However, as in Paper II, the difference is consistent with the strength of system-to-system variation. The SAGA hosts with the 25% most concentrated satellite distributions (as measured by the median projected distance of satellites) have, on average, a satellite radial distribution that is just as concentrated as that of the MW. It is well known that the MW has a more concentrated radial profile when compared with simulations (e.g., Willman et al., 2004; Yniguez et al., 2014), likely due to a combination of the proximity of LMC and SMC to the MW center (Nadler et al., 2020), a more concentrated dark matter distribution due to earlier halo formation, and potential incompleteness in satellite census in the outskirt of the MW. The last reason is unlikely to be an issue when we compare the MW and the SAGA systems given the SAGA mass limit, but both the LMC/SMC locations and the MW’s dark matter halo can contribute to the difference we observe.
@@ 5.6 Comparing Satellite Abundance with Simulations
Overall, the new UM-SAGA model predicts a satellite abundance distribution that agrees with SAGA observations for satellites with . We have conducted two-sample KS-tests using the 100 realizations of 101 SAGA-like hosts versus the actual SAGA data for their distributions, and we found that 94 out of 100 realizations have values and all realizations have , meaning that the UM-SAGA model predictions are statistically indistinguishable from SAGA in terms of distributions. This is consistent with our findings in Paper II, where we compared SAGA Stage II observations to the predictions of the Nadler et al. (2019b, 2020) subhalo abundance matching model. Despite the different empirical assumptions underlying the two predictions, we find that these models yield very similar predictions for SAGA satellite counts. A more detailed comparison between these predictions may inform modeling degeneracies, e.g., related to satellite disruption due to baryonic effects and modeling treatment of orphan galaxies in the simulation.
@@ 5.7 Correlation between Satellite Abundance and System Properties
Photometric properties of the host galaxies: -band absolute magnitude , color, surface brightness, Sersic index, ratio of semi-minor to semi-major axes , position angle;
@@
Other measured or inferred properties of the host galaxies: stellar mass, group mass (Lim et al., 2017, see Section 2.2 for a brief description of how the halo mass estimate was obtained), star formation rate and specific star formation rate from NUV, HI mass, distance to the nearest galaxies with ; and
@@
Given this long list of properties, the look-elsewhere effect can impact the interpretation of the statistical significance of any observed correlation. While we report the measured correlation strengths below, we limit our discussion to a few most significant cases, where -value is below or of the order of . Note that for 101 systems, a -value of 0.046 (2) corresponds to a Spearman correlation coefficient of 0.2 and a -value of (3) corresponds to a coefficient of 0.3.
@@
Figure 15 summarizes the results of our correlation calculation. Each row shows the correlation strength (right is stronger) between the property listed on the left and the number of satellites in the Gold (gold squares) or Gold+Silver (silver circles) samples. The sign on the leftmost edge shows the sense of the correlation. We express the correlation strength in terms of -value because the number of systems that enter the calculation in each row can be different, and the -values already take the number of systems into account.
@@ 5.7.1 Notable strong correlations
Among the properties that we tested, the stellar mass of the most massive satellite has the strongest correlation with satellite abundance, and the correlation strength is significantly stronger than that of any other properties, including the halo mass (obtained from Lim et al. 2017 group catalog; see Section 2.2). One might expect that the halo mass would correlate most strongly with satellite abundance; however, gravity-only simulations have shown that even at the fixed halo mass, satellite abundance will vary with halo properties such as halo formation time (e.g., Mao et al., 2015). On the other hand, if the satellites are simply random draws of some underlying stellar mass function (see e.g., the Poisson point process model proposed by Mao et al., 2015), then both the most massive satellite mass and the satellite abundance at a fixed mass limit are directly dictated by the overall amplitude of the satellite mass function, resulting in a strong correlation between those two quantities.
This effect can be seen most clearly in Figure 16, where colors of the points indicate satellite abundance, and hence the direction of the color gradient tells us how the satellite abundance depends on the two quantities plotted, group mass and most massive satellite mass. We can see that the color gradient goes vertically, supporting the hypothesis that the overall amplitude of the satellite mass function drives the correlation between satellite abundance and the most massive satellite mass. A similar correlation is also observed in nearby MW-mass systems (Smercina et al., 2022). Figure 16 also shows that at a fixed group (halo) mass, the amplitude of the satellite mass function still varies, likely with other halo properties that we are not able to measure here.
@@
We plan to explore these correlations in more detail in follow-up work where we more carefully take into account the inter-correlations among the different properties, which will allow us to distinguish the primary correlation. In addition, in Appendix A, we also show how the satellite SMF as a whole changes with a selection of the system properties presented here.

### v3

@@ 3.4 Combining Spectroscopic and Photometric Data Sets
As the photometric catalog has been updated a few times during the course of the SAGA Survey, we combine the spectroscopic and photometric measurements by matching the sky coordinates. For each spectroscopic measurement, we search the photometric catalog within a 20′′ radius. Given the SAGA Survey depth, the photometric sources are not too crowded, and in most cases, only one unique match exists. When multiple photometric sources are possible matches, we identify the best match based on the luminosities and sizes of those photometric objects. The specific order by which a photometric source is selected as the best match to a spectroscopic measurement is described in the list in the SAGA code base.
@@ 4.2 Derived Properties
Radius and surface brightness: For most galaxies, the circularized half-light radius is derived directly from the axis ratio and semi-major axis from the Legacy Surveys catalog measured in the band. However, for galaxies available in the Siena Galaxy Atlas, we use as the half-light radius (one-third of the isophotal radius at 26 mag arcsec-2). We also use GALFIT to re-fit the photometry for a handful of galaxies that clearly have a bad photometric fit in the original catalog by visual inspection. We then calculated the effective surface brightness with the following formula:
@@
Stellar mass: As in Paper II, we estimate stellar mass using the Bell et al. (2003) relation, but calibrated the relation to recent estimates from Zibetti et al. (2009) and Taylor et al. (2011). We also assume a Kroupa (2001) initial mass function and an absolute solar -band magnitude of 4.65 (Willmer, 2018). The resulting stellar mass conversion is
@@
For SAGA satellite galaxies (Section 4.3), we also measure the following quantities:
@@
Star Formation Rate (SFR): we calculate the star formation rate based on the both H equivalent width () and near-UV (NUV) flux. See Section 2.4 of Paper IV for details.
@@
“Quenched” flag: we label a galaxy as quenched if it has no significant H emission () and if it has a specific SFR (sSFR) in NUV below . See Section 3.1 of Paper IV for more details.
@@ 4.3 SAGA Satellites
As in Paper II, we define a “satellite” as a galaxy that satisfies the following criteria: (i) fainter than the host galaxy, (ii) within 300 kpc in projection to the host galaxy, and (iii) has a heliocentric velocity that is within 275 km s-1 of the host galaxy. From the SAGA redshift catalog around 101 hosts, we identified 378 satellites in total, with 229 whose redshifts were first obtained by the SAGA Survey. Galaxies that do not meet these criteria are considered to be “background” galaxies (a handful of these galaxies are actually foreground galaxies; however, we will use the term background galaxies to refer to any non-satellites).
@@
Figure 5 demonstrates this definition, with the satellite galaxies being in the middle left part of the plot. It is important to note that this observationally motivated satellite definition is not an attempt to strictly select objects that lie within the three-dimensional physical virial radius of their dark matter host halo, nor to strictly select objects that are gravitationally bound to the host. With only line-of-sight velocities and sky positions available, we chose a definition that is straightforward to implement in both observational and simulation data.
@@
With simulations, we can estimate the fraction of interlopers (galaxies that sit outside the three-dimensional radius in the line-of-sight direction) and potential fly-by galaxies (galaxies that sit within the three-dimensional radius but have a speed greater than the escape velocity; Sinha & Holley-Bockelmann, 2012; An et al., 2019). Based on gravity-only simulations, the SAGA satellite definition gives an average interloper fraction of 30% for MW-mass halos. This fraction varies with projected distance, and drops to 15% for galaxies within 150 kpc in projection, as demonstrated in Paper II. Note that the radius considered here is always fixed to be 300 kpc, even though the individual SAGA systems could have different virial radii, which are not measurable observationally.
@@
As for potential fly-by galaxies, Figure 5 highlights the fact that a small fraction (4%) of SAGA satellites may be unbound, indicated by purple squares (satellites) sitting outside the red solid bands, which represent the circular velocity of an Navarro–Frenk–White (NFW; Navarro et al. 1996) profile with and (red solid lines) plus the typical spectroscopic velocity error of 60 km s-1 in our measurement. The fraction of unbound objects observed is consistent with the gravity-only simulations when we use the SAGA satellite definition to select the halos in the simulations.
@@
Figure 5 also shows the velocity dispersion measured from all SAGA satellites. Satellites are binned in projected distances. For each bin, we fit a single Gaussian distribution, with its mean and variance being the free parameters, convolved with a constant velocity measurement error of 60 km s-1. Because we use all satellites to measure the velocity dispersion profile, it represents the stacked profile where systems with more satellites are weighted more. The velocity dispersion declines with the projected radius to the virial radius, suggesting we are measuring a true dynamical mass. The increase beyond 300 kpc is likely due to increased contribution from interlopers. We are not able to measure the velocity dispersion for individual systems due to the small number of satellites in each system.
@@ 4.4 SAGA Background Sample
This sample of low-redshift and low-mass background galaxies, although not the main focus of the SAGA Survey, has enabled a range of studies, including optimizing machine learning algorithms to select low-mass galaxies (Wu et al., 2022; Darragh-Ford et al., 2023), calibrating photometric sources for low-mass galaxy weak lensing measurements (Thornton et al., 2023), and measuring the mass loading factor in low-mass galaxies (Kado-Fong et al., 2024).
@@ 4.5 Likely Satellite Candidates and Incompleteness Correction
The quenched labels for unconfirmed candidates are determined by the following color–magnitude relation:
@@ 5.1 Science Samples of SAGA Satellites
These three science samples and the numbers of confirmed and potential satellites in each sample are summarized in Table 1. The cutout images of the 378 confirmed satellites in all three samples are shown in Figure 7.
@@ 5.2.1 Individual SAGA stellar mass functions
We first show the cumulative satellite stellar mass function (SMF) for each of the 101 SAGA systems in Figure 8. While it can be difficult to discern individual SMFs in this figure, the main purpose of this figure is to demonstrate the system-to-system spread of the SMFs. The inset of this figure also allows us to quickly see the distribution of the number of satellites for the Gold and GoldSilver samples. Note that the SMFs and the number of satellites shown here already include the incompleteness correction described in Section 4.5 and Section B.1. Four of the 101 systems have the same level of survey completeness but do not have any confirmed satellites; their incompleteness-corrected SMFs are shown in this figure as well.
@@ 5.2.2 Comparison with MW and M31 SMFs
The satellite SMF of MW and M 31 are also shown in Figure 8, and they are both well within the system-to-system spread of the SAGA SMFs. In particular, the existence of an LMC-mass satellite is not uncommon; about one-third (34/101) of the SAGA systems have at least one confirmed satellite with a stellar mass above . In the inset, one can see that the number of satellites in M 31 is on the high end of the SAGA systems, which is expected given M 31’s higher mass compared to SAGA hosts (Figure 3). On the other hand, the number of satellites in the MW is just slightly below the median number of the SAGA systems for the GoldSilver sample. If one only considers the expected correlation between the host galaxy’s stellar mass and the number of satellites, the MW would appear to have a lower satellite number, given that the majority of SAGA hosts are less massive than the MW.
Note that in the main panel, the low-mass end of the MW SMF is closer to the blue-colored SAGA SMFs, while the high-mass end of the MW SMF is closer to the yellow-colored SAGA SMFs. In other words, the number of satellites in the MW (to the SAGA mass limit; i.e., GoldSilver sample) is lower than that of the SAGA systems that also have massive satellites like the LMC. This observation suggests that the MW satellites do not follow the strong correlation between the number of satellites and the most massive satellite mass seen in SAGA; we will demonstrate this correlation quantitatively in Section 5.7. This discrepancy suggests that the MW may have an atypical formation history compared to the SAGA hosts, which we will discuss in detail in Section 6.1.
@@ 5.3 Satellite Quenched Fraction
As Figure 9 demonstrates, we expect that a subset of low-mass quenched satellites exist in these systems for which we do not have a confirmed redshift. Since we have a good estimate of the expected number of these unconfirmed satellites, we include them in the calculation of the quenched fraction. For unconfirmed satellites, the quenched definition is based on their color and absolute magnitude, as described in Eq. 6. Note that the incompleteness-corrected and uncorrected results are very similar in the Gold sample (right of the vertical gold dashed line), and deviate from each other at lower masses.
The satellite quenched fraction in the 101 SAGA systems is marginally higher than, but statistically consistent with, what we reported in our Stage II results (Paper II). When compared with the satellite quenched fraction in the Local Group, SAGA systems’ average satellite quenched fraction is lower for stellar mass below . This difference between SAGA systems and the Local Group is unlikely to be due to observational incompleteness for three reasons:
@@
First, if the SAGA Gold sample had an overall satellite quenched fraction of 70%, like the Local Group, it would imply there were more than 200 quenched satellites with that were not identified by our survey. This is extremely unlikely given our completeness estimates, and would significantly alter the stellar mass function if it were true.
@@ 5.4 Satellite Properties
The color–mass relation (left panel) shows a visible separation between red quenched satellites and blue star-forming satellites, although some star-forming satellites have redder colors than expected for their mass. As a result, there is no simple color–mass cut that can cleanly separate quenched and star-forming satellites. The color–mass cut used by Carlsten et al. (2022) to separate quenched and star-forming satellites is shown as a thin brown line on the plot. Notably, the LMC and the SMC (the two pink stars with the highest stellar masses) are bluer than any of the SAGA satellites of similar mass, possibly as a result of star formation triggered by their ongoing infall into the MW (Shao et al., 2018). We refer readers to Section 7.2 of Paper IV for further discussion.
@@ 5.5 Satellite Radial Distribution
However, as in Paper II, the difference is consistent with the strength of system-to-system variation. The SAGA hosts with the 25% most concentrated satellite distributions (as measured by the median projected distance of satellites) have, on average, a satellite radial distribution that is just as concentrated as that of the MW. It is well known that the MW has a more concentrated radial profile when compared with simulations (e.g., Willman et al., 2004; Yniguez et al., 2014), likely due to a combination of the proximity of LMC and SMC to the MW center (Nadler et al., 2020), a more concentrated dark matter distribution due to earlier halo formation, and potential incompleteness in satellite census in the outskirts of the MW. The last reason is unlikely to be an issue when we compare the MW and the SAGA systems given the SAGA mass limit, but both the LMC/SMC locations and the MW’s dark matter halo can contribute to the difference we observe.
@@ 5.6 Comparing Satellite Abundance with Simulations
Overall, the new UM-SAGA model predicts a satellite abundance distribution that agrees with SAGA observations for satellites with . We have conducted two-sample Kolmogorov–Smirnov tests using the 100 realizations of 101 SAGA-like hosts versus the actual SAGA data for their distributions, and we found that 94 out of 100 realizations have -values and all realizations have , meaning that the UM-SAGA model predictions are statistically indistinguishable from SAGA in terms of distributions. This is consistent with our findings in Paper II, where we compared SAGA Stage II observations to the predictions of the Nadler et al. (2019b, 2020) subhalo abundance matching model. Despite the different empirical assumptions underlying the two predictions, we find that these models yield very similar predictions for SAGA satellite counts. A more detailed comparison between these predictions may inform modeling degeneracies, e.g., related to satellite disruption due to baryonic effects and modeling treatment of orphan galaxies in the simulation.
@@ 5.7 Correlation between Satellite Abundance and System Properties
Photometric properties of the host galaxies: -band absolute magnitude , color, surface brightness, Sersic index, the ratio of semi-minor to semi-major axes , and position angle.
@@
Other measured or inferred properties of the host galaxies: Stellar mass, group mass (Lim et al., 2017, see Section 2.2 for a brief description of how the halo mass estimate was obtained), star formation rate and specific star formation rate from NUV, HI mass, distance to the nearest galaxies with .
@@
Given this long list of properties, the look-elsewhere effect can impact the interpretation of the statistical significance of any observed correlation. While we report the measured correlation strengths below, we limit our discussion to a few most significant cases, where the -value is below or of the order of . Note that for 101 systems, a -value of 0.046 (2) corresponds to a Spearman correlation coefficient of 0.2 and a -value of (3) corresponds to a coefficient of 0.3.
@@
Figure 15 summarizes the results of our correlation calculation. Each row shows the correlation strength (right is stronger) between the property listed on the left and the number of satellites in the Gold (orange-filled squares) or Gold+Silver (gray-filled circles) samples. The sign on the leftmost edge shows the sense of the correlation. We express the correlation strength in terms of -value because the number of systems that enter the calculation in each row can be different, and the -values already take the number of systems into account.
@@ 5.7.1 Notable strong correlations
Among the properties that we tested, the stellar mass of the most massive satellite has the strongest correlation with satellite abundance, and the correlation strength is significantly stronger than that of any other properties, including the halo mass (obtained from the Lim et al. 2017 group catalog; see Section 2.2). One might expect that the halo mass would correlate most strongly with satellite abundance; however, gravity-only simulations have shown that even at the fixed halo mass, satellite abundance will vary with halo properties such as halo formation time (e.g., Mao et al., 2015). On the other hand, if the satellites are simply random draws of some underlying stellar mass function (see, e.g., the Poisson point process model proposed by Mao et al., 2015), then both the most massive satellite mass and the satellite abundance at a fixed mass limit are directly dictated by the overall amplitude of the satellite mass function, resulting in a strong correlation between those two quantities.
This effect can be seen most clearly in Figure 16, where colors of the points indicate satellite abundance, and hence, the direction of the color gradient tells us how the satellite abundance depends on the two quantities plotted, group mass and most massive satellite mass. We can see that the color gradient goes vertically, supporting the hypothesis that the overall amplitude of the satellite mass function drives the correlation between satellite abundance and the most massive satellite mass. A similar correlation is also observed in nearby MW-mass systems (Smercina et al., 2022). Figure 16 also shows that at a fixed group (halo) mass, the amplitude of the satellite mass function still varies, likely with other halo properties that we are not able to measure here.
@@
We plan to explore these correlations in more detail in follow-up work, where we more carefully take into account the inter-correlations among the different properties, which will allow us to distinguish the primary correlation. In addition, in Appendix A, we also show how the satellite SMF as a whole changes with a selection of the system properties presented here.